OpenStudy (anonymous):

What are the roots of 2x+6=−5x/x? 3−+i2 −3−+i2 −3−+2i4 3−+i2√4

OpenStudy (anonymous):

I'll give medal!!!!!!1

OpenStudy (freckles):

I do not see how you can possibly have two answers to this question.

OpenStudy (freckles):

wait now i can't tell what the question is was it 2x+6=5x or something like that

OpenStudy (anonymous):

What are the roots of 2x+6=−5x/x? 3−+i2 −3−+i2 −3−+2i4 3−+i2√4

OpenStudy (freckles):

It keeps showing a bunch of questions

OpenStudy (anonymous):

idk why its doing that

OpenStudy (freckles):

5x/x is just 5

OpenStudy (anonymous):

ill message you it

OpenStudy (freckles):

so you have 2x+6=-5 ?

OpenStudy (freckles):

@stellawalkerson you do not have an x^2 somewhere in the problem?

OpenStudy (tkhunny):

2x+6=−5x/x \(x \ne 0\) \(2x^{2} + 6x = -5x\) \(2x^{2} + 11x = 0\) \(x(2x+11) = 0\) \(x = 0\;or\;x = -11/2\) Thus, \(x = -11/2\) Doesn't matter how you do it or look at it. Still only one.

OpenStudy (anonymous):

Thats not how you do it

OpenStudy (anonymous):

@tkhunny check ur message

OpenStudy (freckles):

you sent me 2x+6=-5x

OpenStudy (freckles):

2x+6=-5x is a linear equation

OpenStudy (tkhunny):

@stellawalkerson Here's the plan. Listen first. Okay, you've done that. Now, show your work. Perhaps you can explain what the actual problem statement is. Then, show your work. As it is posted, that IS how you do it.

OpenStudy (freckles):

a linear equation will never have a 2 root possibility

OpenStudy (freckles):

did you mean 2x+6=-5/x ?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

ok well -5x/x is definitely not the same as -5/x

OpenStudy (freckles):

because -5x/x is actually -5 (assuming the x does not equal 0)

OpenStudy (freckles):

and -5/x is not -5

OpenStudy (freckles):

but anyways tk actually solved 2x+6=-5x/x in a similar way you should solve 2x+6=-5/x

OpenStudy (freckles):

The first step was to multiply x on both sides You will get something you recognize more when you do that

OpenStudy (tkhunny):

@stellawalkerson You still haven't shown ANY work of your own. Do better.